OpenStudy (anonymous):

Salma, Renata, and Rachel will share the cost to rent a vacation house for a week. Salma will pay 40% of the cost. Renata will pay 0.35 of the cost. Rachel will pay the remainder of the cost. A) Rachel thinks that she will pay 1/3 of the cost. Is Rachel correct? B) The cost to rent a vacation house for a week is $850. How much will Salma, Renata, and Rachel each pay to rent this house for a week?

OpenStudy (anonymous):

Okay so 40% of 850 is 340 to make 0.35 a percent u just move the decimals 2 to the right and get 35% 35% of 850 is 297.50 so 340+297.50= 637.50 so 1/3 as a percent is 0.33 again to make that a percent we move the decimals shich gets 33% 33% of 850 is 280.50 637.50+ 280.50 =918 which is more than the amount to rent so A. NO, SHE WILL PAY LESS

OpenStudy (anonymous):

what about part B?

OpenStudy (greenbannanabob):

First change all the values into decimals. Salma-0.4 Renata-0.35 Rachel-0.25 Multiply each value by 850. Salma-0.4*850= Renata-0.35*850= Rachel-0.25*850= The value you get for each one is your answer for B.

OpenStudy (anonymous):

part b is easy too you just have to figure out how much is she actually going to pay which is 25% because of one person is paying 40% and one is paying 35% which 75% of the cost so 25% is missing that is what i think

OpenStudy (anonymous):

thank you!

OpenStudy (greenbannanabob):

You're welcome!